OpenStudy (anonymous):

Graph using intercepts 7x + 2y = 6 I got (0,3) and (6/7, 0) how exactly do I graph the second intercept pair (6/7, 0)?

OpenStudy (3psilon):

Graph the second what?

OpenStudy (anonymous):

(6/7, 0)

OpenStudy (anonymous):

Anyone? I'm really stuck and have a lot of work to finish. I'd really appreciate some help.

hero (hero):

7x + 2y = 6 7(0) + 2y = 6 2y = 6 y = 6/2 y = 3 Therefore, first point is (0,3) 7x + 2(0) = 6 7x = 6 x = 6/7 Therefore, the second point is (6/7, 0) Graph both points on an xy, draw a line through them, then label the line with the given equation.

OpenStudy (anonymous):

But the second point isn't an actual number, it's a fraction. How do I graph (6/7, 0)?

hero (hero):

Hint: 6/7 is close to 1